OpenStudy (anonymous):

Write each of the following in terms of sin θ and cos θ, and then simplify if possible. sec θ cot θ

OpenStudy (anonymous):

\[\sec \theta . \cot \theta = 1/\sin \theta. \cos \theta/\sin \theta\] \[=\cos \theta/\sin ^{2} \theta\]